OpenStudy (tylermckinney16):

Nick has budgeted $60.00 per week for traveling expenses. He will spend $6.00 round trip per day to get to and from school. Write an equation to represent this problem (5 points) and use your equation to find out how much money he will have left after 5 days (5 points).

OpenStudy (3mar):

May I help?

OpenStudy (tylermckinney16):

Yes :)

OpenStudy (3mar):

With my pleasure! Any ideas? Have the key?

OpenStudy (tylermckinney16):

to be honest im completely lost lol idk what to do.

OpenStudy (3mar):

I am with you to get the idea In Sha' Allah! Don't worry... "Nick has budgeted $60.00 per week for traveling expenses" So the \(initial \) amount of money that he will spend from it $60.. agree>

OpenStudy (tylermckinney16):

Yeah i agree

OpenStudy (3mar):

That is good.... next step: "He will spend \(\color{red}{$6.00}\) round trip \(\color{red}{per~ day}\) to get to and from school" So 6$ every day is spent.. agree?

OpenStudy (danjs):

let x = day of the week , y=money remaining after x days in the week The idea is, money remaining = initial amount - costs

OpenStudy (tylermckinney16):

Yep i aggree to that to :)

OpenStudy (3mar):

So take it @DanJS

OpenStudy (danjs):

y = $60 - spending spending is $6 per 1 day, so for 2 days it would be 2*6, or 3 days 3*6, x days 6*x ... the cost is the daily price times the number of days used y = $60 - $6*x

OpenStudy (danjs):

Money remaining is the initial 60 take away 6 dollars per x days gone past x-- days into the week y--money ramaining after x days y = 60 - 6*x

OpenStudy (tylermckinney16):

ok give me a second

OpenStudy (tylermckinney16):

Would it be 30?

OpenStudy (danjs):

yeah they want day 5, x=5

OpenStudy (tylermckinney16):

ok

OpenStudy (tylermckinney16):

What next @DanJS

OpenStudy (3mar):

Can I proceed with you? @TylerMckinney16

OpenStudy (tylermckinney16):

yes please.

OpenStudy (3mar):

Thank you. Pleasure is mine! Where were you?

OpenStudy (3mar):

As @Danjs showed... The initial 60 take away 6 dollars per x days gone past x-- days into the week y--money ramaining after x days y = 60 - 6*x

OpenStudy (tylermckinney16):

My computer died before lol

OpenStudy (tylermckinney16):

And ok

OpenStudy (3mar):

and he is alive now?? ;)

OpenStudy (tylermckinney16):

Yes lol

OpenStudy (3mar):

Two requires: - Write an equation to represent this problem and - use your equation to find out how much money he will have left after 5 days

OpenStudy (tylermckinney16):

yep

OpenStudy (3mar):

Very good.. Are you satisfied with the first part? \[\Huge f(\color{red}x)=60-6\color{red}x\]

OpenStudy (tylermckinney16):

Yeah but how to i work the problem lol?

OpenStudy (3mar):

"how to i work the problem" you mean the second part?

OpenStudy (tylermckinney16):

Yeah

OpenStudy (3mar):

It is easier that the first one! "use your equation to find out how much money he will have \(\color{red}{left}\) after \(\color{red}5\) days " means: how much money remains after 5 days of spending money???

OpenStudy (tylermckinney16):

Would it be 30?

OpenStudy (3mar):

If @Danjs let me quote, "The idea is, money remaining \(f(x)\) = initial amount \(60\) - what has been spent \(6x\)" i.e: \[\Huge f(\color{blue}x)=60-6\color{blue}x\] \[\Huge f(\color{blue}5)=60-6(\color{blue}{5})=~~?\]

OpenStudy (tylermckinney16):

55?

OpenStudy (3mar):

YES... it would be \(30\)..

OpenStudy (3mar):

\[\huge f(\color{blue}5)=60-6(\color{blue}{5})=60-30=30 \checkmark\]

OpenStudy (tylermckinney16):

Oh so thats the answer :)

OpenStudy (3mar):

Exactly... BUT from brother to brother... Stop 2nd guessing yourself! You hit it firstly then retracted...no...trust your answer as you got it with steps..

OpenStudy (tylermckinney16):

Ok :)

OpenStudy (3mar):

You are hard working student and body builder also! ;)